Loretta's age in 11 years will be three times what it was 7 years ago. How old is Loretta now?
Answer by stanbon(75887)

Loretta's age in 11 years will be three times what it was 7 years ago. How old is Loretta now?
-----
Equation:
L+11 = 3(L-7)
---
L + 11 = 3L - 21
----
2L = 32
---
Loretta is 16 yrs. old now
=============================
Cheers,
Stan H.
====================